OpenStudy (anonymous):

4a = 6 - b 4a = 5 - b A. There is one solution. The solution is ____. B. There are infinitely many solutions. C. There are no solutions.

OpenStudy (zzr0ck3r):

multiply the second one by -1 we have 4a = 6-b -4a = -5+b add them 0 = 1 thus no solutions

hartnn (hartnn):

as 4a cannot be simultaneously equal to 6-b and 5-b this system has no solution.

OpenStudy (zzr0ck3r):

that to lol

OpenStudy (anonymous):

I would use substitution to solve this system. Solve for a in the first equation, substitute into the 2nd equation, solve for b. Then check answers.

OpenStudy (zzr0ck3r):

elimination is much faster here

OpenStudy (zzr0ck3r):

but hartn gave the best answer imo

OpenStudy (anonymous):

It said to solve it using substitution... I guessed it was no solution but I wasnt sure.

hartnn (hartnn):

now u are.

OpenStudy (lgbasallote):

as the old people say, "You can't guess and be sure at the same time"

OpenStudy (anonymous):

thank you all!

OpenStudy (lgbasallote):

hah i got thanked for being a smartmouth

OpenStudy (zzr0ck3r):

by substitutaion 4a = 6 - b 4a = 5 - b so a = (5-b)/4 plug this into the first equation 4a = 6-b 4(5-b)/4 = 6-b 5-b = 6-b 5=6 thus no solutions

OpenStudy (zzr0ck3r):

lol @lgbasallote

OpenStudy (lgbasallote):

you could have just substituted 4a to 4a...

OpenStudy (zzr0ck3r):

its late

OpenStudy (anonymous):

yes, I know but i needed to get this done!

Parth (parthkohli):

You could also equate both RHSs. For example,\[y = 3x \\ y = 3 + x \\ \\ \implies 3 + x = 3x\]